# Morpho-Photometric Classification of KiDS DR5 Sources Based on Neural Networks: A Comprehensive Star-Quasar-Galaxy Catalog

### v1

@@ 4.4 Model Interpretation
Neural network models are often regarded as “black boxes” with internal decision-making processes that are opaque to humans. This opacity makes it difficult to directly analyze the intrinsic connections between celestial object classification and their physical properties. Here, we attempt to use empirical knowledge to explain the features that the model may have learned during the its training, aiming to better understand the physical mechanisms behind its excellent performance.
@@
The proposed multimodal neural network extracts information from two input branches: SED and image. As shown in Figure 11, different types of celestial objects inherently exhibit distinct SED and morphological characteristics. The SED primarily reflects the differences in internal radiation mechanisms and the distances (redshifts) of the objects. Stars, for instance, typically exhibit SEDs that can be approximated by blackbody radiation, with peak positions mainly determined by their effective temperatures. On the other hand, the spectrum of quasars in the ultraviolet to optical bands is dominated by the “big blue bump”, a feature produced by high-temperature radiation from the accretion disk, while the near-infrared is predominantly emission from hot dust (Sanders et al., 1989; Elvis et al., 1994). The multi-wavelength luminosity of galaxies is more complex, as it involves the combined effects of stellar populations and dust (Conroy, 2013). The nine-band photometric data provided by KiDS spans a broader wavelength coverage than SDSS spectra, potentially yielding additional physical information. For example, the SED peak of M-type stars is located in the near-infrared band (see Figure 11), and as the redshift increases, the contribution of the “big blue bump” in quasars becomes more pronounced in the optical-near-infrared bands. This prior knowledge suggests that the model may achieve classification by learning the relative brightness across multiple bands.
@@
For the imaging side, it can provide rich details of the morphology of celestial objects. These morphological features not only contain information about the physical nature of objects but might also refine flux measurements, as the shapes of different objects can affect the results of GAaP measurements. To test the impact of images on classification, we separately used MLP to classify the SED (see Appendix), and found that the performance was significantly inferior to the multimodal neural network. This demonstrates that incorporating morphological information can indeed significantly improve classification efficacy. To further confirm which information directly aids classification, we normalized all images and found that there was a negligible effect on model performance. This implies that during the classification process, the model primarily relies on the physical morphological features rather than simply re-measuring the flux from the images. This is consistent with the human inferential strategies, where we prefer to use the physical properties of morphology itself rather than mere flux correction when classifying. For example, extended galaxies are easily distinguishable from stellar point sources, and the size can be used to estimate the distance of the galaxy (see also Figure 9). For high-redshift quasars and compact galaxies, they may appear more like stars in images, but we can at least rule out the possibility of them being low-redshift extragalactic objects.
@@
In summary, the multimodal neural network can extract luminosity information from photometric data and morphological features from image data, and mapping both to a high-dimensional feature space via a nonlinear function. Within this space, various categories are effectively separated, thus achieving the purpose of classification. The accuracy and efficiency of this method hold promise for widespread application in next-generation multi-color survey projects such as LSST, WFST (Wang et al., 2023), and Mephisto (Han et al., 2022). Future space survey projects (e.g., CSST and Euclid) will provide unprecedented high-resolution images and slitless spectral data, which will further refine our model and enhance the precision of astronomical classification. Moreover, the anticipated wealth of spectral data from future spectroscopic surveys will enable our model to extend its performance to fainter magnitudes.
@@ 5 Conclusion
In this work, we have developed a multimodal neural network architecture that effectively combines morphological information from images and multi-band photometry for the classification of stars, quasars, and galaxies in the KiDS DR5 survey. By training on SDSS spectroscopic labels and testing on both internal and external datasets, we demonstrate the model’s high accuracy, robustness, and ability to generalize to new data. Our key findings are:
@@
This work is supported by National Key R&D Program of China (No. 2021YFA1600404), the National Natural Science Foundation of China (NSFC grants No. 12303022, 12203096, 12203050, 12373018, 11991051, and 12203041), Yunnan Fundamental Research Projects (grants NO. 202301AT070358 and 202301AT070339), Yunnan Postdoctoral Research Foundation Funding Project, Special Research Assistant Funding Project of Chinese Academy of Sciences, and the science research grants from the China Manned Space Project with No. CMS-CSST-2021-A06, NRN acknowledges financial support from the NSFC, Research Fund for Excellent International Scholars (grant No. 12150710511), and from the research grant from China Manned Space Project n. CMS-CSST-2021-A01.
@@ Appendix A Performance of MLP
To assess the contribution of morphological information in our multimodal neural network, we conducted an independent performance evaluation of the MLP branch. In this test, the MLP branch used only the SED features from the KiDS dataset, which include photometry from 9-bands and 8 adjacent band colors, without integrating the morphological features processed by the CNN branch. The MLP was trained and evaluated using the same training and testing datasets as the full multimodal network. Figure 12 shows the evaluation metrics for the MLP at different output probability thresholds, analogous to Figure 3 for the full multimodal network.
@@
The MLP achieves an overall accuracy of 97.15% on the testing dataset, with F1 scores of 96.13%, 89.90%, and 98.49% for stars, quasars, and galaxies at the default threshold. While still good, these metrics are noticeably lower than the performance of the multimodal network, especially for quasars. The MLP maintains high purity for galaxies and stars as the threshold is raised, but suffers a faster drop in completeness compared to the multimodal network. For quasars, the MLP fails to achieve both high purity and high completeness at any threshold, whereas the multimodal network can reach 99% purity with 83% completeness at $p\geqslant 0.99$.
@@
These results demonstrate that the morphological information extracted by the CNN branch contributes significantly to the high performance of the multimodal network, especially when dealing with objects that are difficult to distinguish morphologically. In this way, the multimodal neural network effectively leverages complementary information from different data sources to achieve high-accuracy classification of stars, quasars, and galaxies.

### v2

@@ 5.1 Model Interpretation
Neural network models are often regarded as “black boxes” with internal decision-making processes that are opaque to humans. This opacity makes it difficult to directly analyze the intrinsic connections between celestial object classification and their physical properties. Here, we attempt to use empirical knowledge to explain the features that the model may have learned during its training, aiming to better understand the physical mechanisms behind its excellent performance.
@@
The proposed MNN extracts information from two input branches: SED and image. As shown in Figure 10, different types of celestial objects inherently exhibit distinct SED and morphological characteristics. The SED primarily reflects the differences in internal radiation mechanisms and the distances (redshifts) of the objects. Stars, for instance, typically exhibit SEDs that can be approximated by blackbody radiation, with peak positions mainly determined by their effective temperatures. On the other hand, the spectrum of quasars in the ultraviolet to optical bands is dominated by the “big blue bump”, a feature produced by high-temperature radiation from the accretion disk, while the near-infrared is predominantly emission from hot dust (Sanders et al., 1989; Elvis et al., 1994). The multi-wavelength luminosity of galaxies is more complex, as it involves the combined effects of stellar populations and dust (Conroy, 2013). The nine-band photometric data provided by KiDS spans a broader wavelength coverage than SDSS spectra, potentially yielding additional physical information. For example, the SED peak of M-type stars is located in the near-infrared band (see Figure 10), and as the redshift increases, the contribution of the “big blue bump” in quasars becomes more pronounced in the optical-near-infrared bands. This prior knowledge suggests that the model may achieve classification by learning the relative brightness across multiple bands.
@@
For the imaging side, it can provide rich details of the morphology of celestial objects. These morphological features not only contain information about the physical nature of objects but might also refine flux measurements, as the shapes of different objects can affect the results of GAaP measurements. To test the impact of images on classification, we separately used MLP to classify the SED (see Appendix A), and found that the performance was significantly inferior to the MNN. This demonstrates that incorporating morphological information can indeed significantly improve classification efficacy. To further confirm which information directly aids classification, we normalized all images and found that there was a negligible effect on model performance. This implies that during the classification process, the CNN branch primarily relies on the physical morphological features rather than simply re-measuring the flux from the images.
Moreover, if images provide more accurate flux measurements, directly utilizing images from all bands should yield more precise SED results. To investigate this, we conducted the following tests using all $ugri$ images released by KiDS (for a fair comparison, we also utilized the corresponding 4-band photometric data):
  1. 1.
    We retrained the MNN shown in Figure 2, but the SED branch only used photometric data from the $ugri$ bands (along with colors from 3 adjacent bands).
  2. 2.
    We removed the SED branch from Figure 2 and added 3 additional CNN branches (resulting in 4 CNN branches in total), with each branch taking image data from a single band as input. The model was then trained directly using images from all $ugri$ bands.
  3. 3.
    Based on the model with 4-branch CNN neural network, we added the SED branch for joint training.
The test results are listed in Table 1. The model trained using images from all four bands performed worse than the MNN, indicating that images may not provide more accurate flux measurements than the photometric data. When considering both SED and images, increasing the number of input images did not lead to significant performance improvements. This suggests that images do not play a crucial role in correcting photometric accuracy, as using the image from the same band should provide more accurate corrections for a specific band. These findings also demonstrates the efficiency and practicality of the multimodal approach, which achieves comparable performance using only a single-band image combined with a few photometric data points. This significantly reduces the computational and storage resources required, compared to models that rely on processing multiple image bands.
This is consistent with the human inferential strategies, where we prefer to use the physical properties of morphology itself rather than mere flux correction when classifying. For example, extended galaxies are easily distinguishable from stellar point sources, and the size can be used to estimate the distance of the galaxy (see also Figure 8). For high-redshift quasars and compact galaxies, they may appear more like stars in images, but we can at least rule out the possibility of them being low-redshift extragalactic objects.
@@
In summary, the MNN can extract luminosity information from photometric data and morphological features from image data, and mapping both to a high-dimensional feature space via a nonlinear function. Within this space, various categories are effectively separated, thus achieving the purpose of classification. The accuracy and efficiency of this method hold promise for widespread application in next-generation multi-color survey projects such as LSST, WFST (Wang et al., 2023), and Mephisto (Han et al., 2022). Future space survey projects (e.g., CSST and Euclid) will provide unprecedented high-resolution images and slitless spectral data, which will further refine our model and enhance the precision of astronomical classification. Moreover, the anticipated wealth of spectral data from future spectroscopic surveys will enable our model to extend its performance to fainter magnitudes.
Figure 11: $gri$ color-composited images of some misclassified targets. The vertical axis represents the spectral classification labels from SDSS, and the horizontal axis represents the predictions made by the MNN.
5.2 Limitations and Challenges
Several studies have shown that tree-based ML methods, such as XGBoost (Chen & Guestrin, 2016) or Random Forest (Breiman, 2001), often outperform other algorithms when performing classification tasks on structured data like photometry and colors (e.g., Zeraatgari et al., 2024; Nakoneczny et al., 2019). To investigate this, we trained the widely used XGBoost algorithm on the same training data as our MNN. The results indicate that XGBoost indeed performs better than MLP, but falls short of our MNN (Table 1). This can be primarily attributed to the absence of morphological information in the training process. We note that Nakoneczny et al. (2021) achieved comparable performance to our method by incorporating structured morphological information (stellarity index, Bertin & Arnouts, 1996) into the XGBoost model. However, one of our goals is to develop a flexible and versatile approach that can directly extract morphological information from images. This approach not only enables real-time classification of image data but also captures the high-dimensional, non-linear physical properties underlying the data (LeCun et al., 2015). DNNs are particularly well-suited for this task due to advantages such as their ability to automatically extract complex and hierarchical features directly from raw data, as well as their effectiveness in modeling high-dimensional and nonlinear relationships.
In addition to the choice of ML algorithms, data quality fundamentally determines the upper limit of model performance. Clean data allows the model to focus on learning the underlying patterns that are consistent with the real-world problems during the training process, thereby enhancing its stability and interpretability. Conversely, dirty data may mislead the learning process of model parameters and evaluation results, increasing the risk of overfitting. Indeed, there are some misclassified targets in the SDSS spectral data (Bolton et al., 2012; Lyke et al., 2020). However, since the proportion of these misclassifications is relatively low (about only a few percent), the model still performs impressively well on the testing dataset. This can also be attributed to the hierarchical feature extraction of DNNs, which allows them to reduce the impact of random noise during the training process, as well as the use of techniques such as regularization and batch training (Zhang et al., 2016; Goodfellow et al., 2016). Through random inspection of misclassified samples, we found that some SDSS labels themselves were erroneous, and our MNN model was even able to effectively correct these errors, especially for galaxies misclassified as stars by SDSS (as shown in Figure 11).
The performance of the model is also influenced by the data volume. A larger dataset can provide richer and more detailed feature representations, which helps the model learn complex patterns and effectively mitigate noise and outlier. This, in turn, enhances its accuracy and generalization ability. In our training, the number of quasar samples was significantly less than that of stars and galaxies, which somewhat limited the recognition capabilities of the model for these targets. To better understand the impact of training data volume on classification performance, we conducted a test in Appendix B by artificially selecting subsets of the training data. The results show that the performance of all three source types improves as the training data volume increases, with quasars showing the most significant improvement, followed by stars, and galaxies benefiting the least. This suggests that the model’s performance is sensitive to data volume for sources with more complex or less abundant features. Although our method has demonstrated impressive classification performance, there is still room for further improvement as spectral data continues to accumulate in the future.
In our extrapolation tests, we divided the data into different magnitude bins based on the $r$-band magnitude and used the data from fainter bins for testing. The results indicate that the model can extrapolate up to 1 mag fainter than the training data. However, this test has inherent limitations in terms of quantitatively analyzing the data and feature space. For example, enforcing equal magnitude intervals for each bin results in varying numbers of training samples across bins. Conversely, ensuring an equal number of samples in each bin alters the magnitude intervals, which in turn affects the structure of the feature space. Additionally, regardless of these considerations, the relative proportions of stars, QSOs, and galaxies unavoidably vary across bins, and controlling the distribution of color features remains challenging.
More importantly, the high-dimensional feature space utilized by the DNN may rely on latent features that are beyond our explicit control or understanding. Despite these challenges, as shown in Table 2, the performance differences across subsets—whether caused by variations in magnitude intervals or sample sizes—are generally within 0.5% to 4%. Among these, the 17.5–19.5 and 19–20 magnitude bins are particularly notable because they share identical magnitude intervals and have nearly equal sample sizes (43,905 and 42,849, respectively). The performance differences between these two bins closely align with the expected test variations, suggesting that the extrapolated results remain acceptable within these limits.
@@ 6 Conclusion
In this work, we have developed an MNN architecture that effectively combines morphological information from images and multi-band photometry for the classification of stars, quasars, and galaxies in the KiDS DR5 survey. By training on SDSS spectroscopic labels and testing on both internal and external datasets, we demonstrate the model’s high accuracy, robustness, and ability to generalize to new data. Our key findings are:
@@
We thank the anonymous referee for their valuable comments and suggestions, which greatly improved the quality of this work. We also thank Yuxuan Pang for his discussions. This work is supported by National Key R&D Program of China (No. 2021YFA1600404), the National Natural Science Foundation of China (NSFC grants No. 12203096, 12303022, 12203050, 12373018, 11991051, and 12203041), Yunnan Fundamental Research Projects (grants NO. 202301AT070358 and 202301AT070339), Yunnan Postdoctoral Research Foundation Funding Project, Special Research Assistant Funding Project of Chinese Academy of Sciences, and the science research grants from the China Manned Space Project (No. CMS-CSST-2021-A06, CMS-CSST-2025-A02, CMS-CSST-2025-A07), Hunan Outstanding Youth Science Foundation(2024JJ2040), NRN acknowledges financial support from the NSFC, Research Fund for Excellent International Scholars (grant No. 12150710511), and from the research grant from China Manned Space Project No. CMS-CSST-2021-A01.
@@ Appendix A Performance of MLP
To assess the contribution of morphological information in our MNN, we conducted an independent performance evaluation of the MLP branch. In this test, the MLP branch used only the SED features from the KiDS dataset, which include photometry from 9-bands and 8 adjacent band colors, without integrating the morphological features processed by the CNN branch. The MLP was trained and evaluated using the same training and testing datasets as the full multimodal network. Figure 12 shows the evaluation metrics for the MLP at different output probability thresholds, analogous to Figure 3 for the full multimodal network.
@@
The MLP achieves an overall accuracy of 97.15% on the testing dataset, with F1 scores of 97.20%, 90.41%, and 98.40% for stars, quasars, and galaxies at the default threshold (Table 1). While still good, these metrics are noticeably lower than the performance of the multimodal network, especially for quasars. The MLP maintains high purity for galaxies and stars as the threshold is raised, but suffers a faster drop in completeness compared to the multimodal network. For quasars, the MLP fails to achieve both high purity and high completeness at any threshold, whereas the multimodal network can reach 99% purity with 83% completeness at $p\geqslant 0.99$.
@@
These results demonstrate that the morphological information extracted by the CNN branch contributes significantly to the high performance of the multimodal network, especially when dealing with objects that are difficult to distinguish morphologically. In this way, the MNN effectively leverages complementary information from different data sources to achieve high-accuracy classification of stars, quasars, and galaxies.
Appendix B Impact of Training Data Size on DNN Performance
DNNs are widely recognized for their ability to model complex, non-linear relationships in data, but they often require large training datasets to achieve optimal performance. To investigate how the size of the training dataset affects the classification performance of our MNN, we conducted a series of experiments. For each experiment, we randomly selected subsets of the original training data with sizes ranging from 5,000 to 100,000 samples. The same network architecture and hyperparameters, as described in Section 3, were used for all training runs. Each trained model was then evaluated on the same testing dataset detailed in Section 4.1.1 to ensure consistency.
Figure 13 presents the F1-scores for stars, quasars, and galaxies as a function of training data size. The results reveal a clear trend: the performance of the model improves consistently with increasing training data volume for all three source types.
Interestingly, galaxies achieve the highest F1-scores across all training data sizes, even when trained with relatively small datasets. This robustness likely arises from their distinct morphological features, which make them easier to classify. The F1-scores for stars improve steadily with increasing training data size, though the rate of improvement is less pronounced compared to quasars. This suggests that while a moderate amount of data is sufficient to achieve reasonable performance for stars, additional data still provides incremental benefits. Quasars, on the other hand, exhibit the most significant improvement in F1-score as the training dataset grows. This trend indicates that quasars possess more complex or diverse features compared to stars and galaxies, necessitating larger datasets for the network to generalize effectively.
These results underscore the critical role of data volume in training DNNs, particularly for classes with more complex or less distinct features, such as quasars. The performance saturation observed for galaxies, and to a lesser extent for stars, suggests that their feature spaces are relatively simpler for the network to learn, even with smaller training samples. In contrast, the significant improvement in quasar classification with increasing data highlights the importance of expanding the training dataset for these objects in future studies to further enhance model performance.
Figure 13: F1-scores for stars, quasars, and galaxies as a function of training dataset size. The total training dataset size ranges from 5,000 to 100,000. For each training run, the same validation and testing datasets are used to ensure consistency during the training and evaluation processes.